OpenStudy (anonymous):

Find the rate of change of the distance between the origin and a moving point on the graph of y = x²+1.

OpenStudy (anonymous):

Let L be the distance between the origin and a moving point. I know that L = √(x² + y²). How can I find the rate of change of L? Should I derivative it? In respect of what?

OpenStudy (anonymous):

You know that \[y=x^2+1\] so you can have L(x) as the distance from the origin as a function of x...

OpenStudy (anonymous):

presumably it's with respect to time... d/dt (L= sqrt(x^2 + (x^2+1)^2)) or... d/dt(L^2 = x^2 +y^2) 2L* dL/dt = 2x* dx/dt +2y* dy/dt either way works..

OpenStudy (anonymous):